OpenStudy (angelwings996):

let f(x)=3x-6 and g(x)=x-2. Find f/g and its domain.

OpenStudy (aravindg):

f/g=f(x)/g(x)

Parth (parthkohli):

\[\dfrac{f}{g} =\dfrac{3x - 6}{x - 2} = \dfrac{3(x - 2)}{x - 2} = \cdots\]

OpenStudy (aravindg):

cant u do that now?

Parth (parthkohli):

Now, for the domain: Do you know what the domain is?

OpenStudy (angelwings996):

3 ?

Parth (parthkohli):

Yes :)

Parth (parthkohli):

Can you tell what "domain" means?

OpenStudy (angelwings996):

domain is the set of all possible inputs

Parth (parthkohli):

Yes, so can you tell what possib... hmm wait. We can instead state the set of all inputs that aren't allowed!

OpenStudy (angelwings996):

Someone told me that x = 2 is not a domain to this problem, is that right ?

Parth (parthkohli):

Yes

OpenStudy (angelwings996):

So would the answer be 3; all real numbers except x = 2?

Parth (parthkohli):

Jea.

OpenStudy (angelwings996):

Okay, Thank you for your help! (:

Parth (parthkohli):

=)